OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation. 8x - y = -9 4x - 3y = -22

OpenStudy (anonymous):

A. 4(-8x-9)-3y=-22 B. 4x-3(-8x-9)=-22 C. 4(8x+9)-3y=-22 D. 4x-3(8x+9)=-22

OpenStudy (anonymous):

I am completely confuzzled. I have no idea how or what to do... :/

OpenStudy (lucaz):

first solve the 1st equation for y y=9+8x

OpenStudy (lucaz):

the plug in the second equation 4x-3(8x+9)=-22

OpenStudy (anonymous):

Thank you sooo much! I have been trying this forever and got no where. I filled up an entire page working on this problem. Somehow, I wasn't able to get it on my own. It seems fairly simple. This is the beginning of my second semester of Algebra 2 and it has been a while since I completed semester1..

OpenStudy (lucaz):

=] algebra is always simple

OpenStudy (anonymous):

I wish it was for me :P Help on another question? Its a word problem. I suck at those

OpenStudy (lucaz):

what is it?

OpenStudy (anonymous):

Two competing cable companies have recently restructured their payment plans. Clear TV requires a $30 activation fee and charges $75 per month for their basic plan. Link TV requires no activation fee and charges $82.50 per month for their basic plan. When will the two plans be equal in cost? A.They will not ever be equal B.At 2 months C.At 5 months D.At 4 months

OpenStudy (lucaz):

Clear: 30+75(m) Link 0+82.5(m) set Clear = Link and solve for m 30+75m=82.5m m=4 months

OpenStudy (anonymous):

You're the best!